9. The physician ordered heparin 15 units/kg/hr for a patient who suffered an MI. On hand is 25,000 units of Heparin in 500 mL D5W. The patient weighs 157 lbs. What is the IV flow rate in mL/hr?
Answer #1

ANSWER.

* Calculation of ml/hr.

        ml/hr = Volume(ml) / Time(hr).

      Calculation of volume(ml)

         Volume(ml) = Desired dose/Dose in hand * Quantity.

    Here

         Ordered dose = 15 unit/kg/hr.

         Weight of the patient = 157 lbs.

     Convert lbs to kg for calculation

                   1 lbs = 0.45 kg

        Weight of the patient = 157 * 0.454

                                          = 71.3 kg

         Desired dose(D) for this patient = 15 * 71.3

                                                             = 1069.5 units/hr.

               Dose in hand (H) = 25000 units

                   Quantity (Q) = 500 ml

     As per above formula

             Volume(ml) of drug required = D/H * Q

                                                           = 1069.5/25000 * 500

                                                           = 0.043 * 500

                                                            = 21.5 ml

                                      Time(hr) = 1 hr.

              So

                   Flow rate (ml/hr) = 21.5 ml/hr.